In the lab, Austin has two solutions that contain alcohol and is mixing them with each other. Solution A is
40% alcohol and Solution B is 10% alcohol. He uses 800 milliliters of Solution A. How many milliliters of Solution B does he use, if the resulting mixture is a 20% alcohol solution?

Found 2 solutions by ewatrrr, lwsshak3:
Answer by ewatrrr(24785)

.40*800 + .10x = .20(800+x)
.20(800)/.10 = x

Answer by lwsshak3(11628)

In the lab, Austin has two solutions that contain alcohol and is mixing them with each other. Solution A is
40% alcohol and Solution B is 10% alcohol. He uses 800 milliliters of Solution A. How many milliliters of Solution B does he use, if the resulting mixture is a 20% alcohol solution?
***
let x=amt of solution B used
10%x+40%*800=20%(800+x)
.10x+320=160+.20x
.10x=160
x=1600
How many milliliters of Solution B does he use? 1600